100 Bond Price Coupon (paid annually) Time to Maturity 100 7% 5% undated 2 yearsIf your investment horizon is 6 years, how would you construct a portfolio in which interest rate and reinvestment risk are offset?                                                      
Answer #1

Answer

The formula to calculate Yield to Maturity (YTM) is:

YTM = C + (F-P / n) / (F+P) / n

In case of A,

= 0.07 + 100/6/ 100/2

= 0.3347

= 33.47%

In case of B,

=0.05 + 100/2 / 100/2

= 1.001

= 100%

Hence, it can be interpreted that bond A has low risk in comparison to that of bond B as it has the highest YTM based on the above evaluation. If a potential investor wants to offset reinvestment risk then it is better to choose bond A.